Find the general rule of the sequence: 32, -16, 8, -4, ...
Answer by math_tutor2020(3817)


Without more info, there could be infinitely many possible answers. See my previous post here for more details.

If we assume the sequence is geometric, then,
a = 32 = first term
r = -1/2 = common ratio
The idea is to multiply each term by -1/2

which leads to,
nth term = a*( r )^(n-1)
nth term = 32*( -1/2 )^(n-1)
Again this all hinges on the assumption we have a geometric sequence.

Let's check the answer with something like n = 3
a_n = 32*(-1/2)^(n-1)
a_3 = 32*(-1/2)^(3-1)
a_3 = 32*(-1/2)^(2)
a_3 = 32*(1/4)
a_3 = 8
We have 8 as the third term, which indeed matches with the given sequence. I'll let you check the other values.